# JLAB-THY-24-4244

### v1

@@ VI Conclusions
We explored jet substructure observables, focusing on the ECF, $e_{2}^{\alpha}$, and jet angularity, $\lambda^{\alpha}$, measured on heavy-flavor jets in $e^{+}\,e^{-}$ collisions. By calculating their cumulative distributions for both ungroomed and mMDT-groomed cases, we achieved NLL accuracy level, resumming logarithms of the observable, $\log v$, and of the quark mass to hard scale ratio, $\log\xi$. While the two observables share an identical logarithmic structure, they diverge at terms beyond NLL. Fixed-order contributions were also incorporated to better capture the intricate dead-cone transition region.
@@
We also extended our comparisons to the future ILC setup at a center-of-mass energy of $\sqrt{s}=2$ TeV. Here, the smaller dead-cone angle ($\theta_{D}\approx 0.005$) introduces significant challenges for detecting the effect in $b$-quark jets, emphasizing the need for exceptional detector resolution. Consequently, we propose that analyzing archived LEP data offers a more practical avenue for studying $b$-quark jet substructure than relying solely on future high-energy experiments. That said, exploring the dead-cone effect in top-quark jets remains an exciting prospect for next-generation colliders like the ILC and FCC [116].
@@ Acknowledgements.
We express our gratitude to Simone Marzani and Gregory Soyez for encouraging us to write this paper as well as for many fruitful discussions. OF also would like to thank Yang-Ting Chien for useful and interesting discussions. The work of PKD is supported by European Commission MSCA Action COLLINEAR-FRACTURE, Grant Agreement No. 101108573, and by the Spanish Government (Agencia Estatal de Investigación MCIN/AEI/10.13039/501100011033) Grants No. PID2020-114473GB-I00, No. PID2023-146220NB-I00 and No. CEX2023-001292-S. The work of OF is supported in part by the US Department of Energy (DOE) Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab and by the Department of Energy Early Career Award grant DE-SC0023304. The work of AG and is supported by the Italian Ministry of Research (MUR) by ICSC Spoke 2 under grant BOODINI. Most of the simulation is conducted with computing facilities of the Galileo cluster at the Department of Physics and Astronomy of Georgia State University. The generated events were analysed using the R
@@
framework [117, 118]. All plots in this paper were generated with Matplotlib python library [119]. Our analytical predictions and analysis files are available on request.
@@ Appendix A $\mathcal{O}(\alpha_{s})$ results for $\lambda^{\alpha}$ and $e_{2}^{\alpha}$
×$\times\Bigg{\{}\int^{1}_{0}\operatorname{d}\!z\frac{1+(1-z)^{2}}{2z}\log\mathopen{}\mathclose{{}\left(\frac{(1-z)^{\frac{2}{\alpha}}+((1-z)z)^{\frac{2}{\alpha}}x}{1+((1-z)z)^{\frac{2}{\alpha}}x}}\right)$
@@ Appendix B All-order results for light-quark jets
×$\times\exp\mathopen{}\mathclose{{}\left[-2\mathopen{}\mathclose{{}\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,0)-R_{q}^{\text{(f.c.)}}(v,0)}\right)}\right].$ (19)
@@
Note that $\Sigma_{V}^{q}(v,\xi)$ retains a dependence on $\xi$ because the integral over the running coupling is performed in the decoupling scheme, see Eqs. (12a) and (12). The groomed cumulative distribution is built in the same way and it is given by
@@
×$\times\exp\mathopen{}\mathclose{{}\left[-2\mathopen{}\mathclose{{}\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,0)-R_{q}^{\text{(f.c.)}}(v,0)}\right)}\right],$ (22)

### v2

@@ VI Conclusions
We explored jet substructure observables, focusing on the ECF, $e_{2}^{\alpha}$, and jet angularity, $\lambda^{\alpha}$, measured on heavy-flavor jets in $e^{+}\,e^{-}$ collisions. By calculating their cumulative distributions for both ungroomed and mMDT-groomed cases, we achieved NLL accuracy level, resumming logarithms of the observable, $\log v$, and of the quark mass to hard scale ratio, $\log\xi$. While the two observables share an identical logarithmic structure, they diverge at terms beyond NLL. Partial fixed-order contributions were also incorporated to better capture the intricate dead-cone transition region.
@@
We also extended our comparisons to the future ILC setup at a center-of-mass energy of $\sqrt{s}=2$ TeV. Here, the smaller dead-cone angle ($\theta_{D}\approx 0.005$) introduces significant challenges for detecting the effect in $b$-quark jets, emphasizing the need for exceptional detector resolution. Consequently, we propose that analyzing archived LEP data offers a more practical avenue for studying $b$-quark jet substructure than relying solely on future high-energy experiments. That said, exploring the dead-cone effect in multi-jet events and top-quark jets remains an exciting prospect for next-generation colliders like the ILC and FCC [115].
@@ Acknowledgements.
We express our gratitude to Simone Marzani and Gregory Soyez for encouraging us to write this paper as well as for many fruitful discussions. OF also would like to thank Yang-Ting Chien, Max Knobbe, Simon Plätzer and Andreas Papaefstathiou for useful and interesting discussions. The work of PKD is supported by European Commission MSCA Action COLLINEAR-FRACTURE, Grant Agreement No. 101108573, and by the Spanish Government (Agencia Estatal de Investigación MCIN/AEI/10.13039/501100011033) Grants No. PID2020-114473GB-I00, No. PID2023-146220NB-I00 and No. CEX2023-001292-S. The work of OF is supported in part by the US Department of Energy (DOE) Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates Jefferson Lab and by the Department of Energy Early Career Award grant DE-SC0023304. The work of AG is supported by the Italian Ministry of Research (MUR), and by ICSC Spoke 2 under grant BOODINI. Most of the simulation is conducted using the computing facilities of the Galileo cluster at the Department of Physics and Astronomy of Georgia State University. The generated events are analysed using the R
@@
framework [116, 117]. All plots in this paper are generated with the Matplotlib Python library [118]. Our analytical predictions, MC run cards, and analysis files are available on request.
@@ Appendix A $\mathcal{O}(\alpha_{s})$ results for $\lambda^{\alpha}$ and $e_{2}^{\alpha}$
$\times\Bigg{\{}\int^{1}_{0}\operatorname{d}\!z\frac{1+(1-z)^{2}}{2z}\log\mathopen{}\mathclose{{\left(\frac{(1-z)^{\frac{2}{\alpha}}+((1-z)z)^{\frac{2}{\alpha}}x}{1+((1-z)z)^{\frac{2}{\alpha}}x}}}\right)$
@@ Appendix B All-order results for light-quark jets
$\times\exp\mathopen{}\mathclose{{\left[-2\mathopen{}\mathclose{{\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,0)-R_{q}^{\text{(f.c.)}}(v,0)}}\right)}}\right].$ (19)
@@
Note that $\Sigma_{V}^{q}(v,\xi)$ retains a dependence on $\xi$ because the integral over the running coupling is performed in the decoupling scheme, see Eqs. (12a) and (IV). The groomed cumulative distribution is built in the same way and it is given by
@@
$\times\exp\mathopen{}\mathclose{{\left[-2\mathopen{}\mathclose{{\left(\mathcal{R}^{(\text{f.o.})}_{V}(v,0)-R_{q}^{\text{(f.c.)}}(v,0)}}\right)}}\right],$ (22)